water is flowing through a pipe with a constriction. The diameter of the narrow section is one-half the diameter of the wide section. If the velocity of the incompressible

fluid is 3,20 m/s in the wide section, then what is the velocity of the fluid in the narrow section?


1
Expert's answer
2022-06-06T08:54:59-0400
"v_1A_1=v_2A_2"

"v_1d_1^2=v_2d_2^2"

"v_2=v_1d_1^2/d_2^2=3.20\:\rm m/s*2^2=12.8\:\rm m/s"